Determine whether each system is independent, dependent, or
inconsistent.
y=3x-5
y=3x+2

Answer by stanbon(75887)

Determine whether each system is independent, dependent, or
inconsistent.
y=3x-5
y=3x+2
---------------
The two equations have the same slope
and different y-intercepts so the
system is inconsistent. They have
no point in common.
=====================================
Algebraic demonstration:
Substitute to get:
3x-5 = 3x=2
-5 = 2
That is a contradiction.
The system is inconsistent.
==========================================
cheers,
Stan H.